# ToDo (#1): #2

### v3

@@ 3.4.4 Characterization of Spiral Inflector
The result of adding optimized focusing quadrupoles on the beam distribution entering the spiral inflector is illustrated in Fig. 3.16.
Figure 3.16: Illustration of beam tightness at the entrance of the spiral inflector before and after the addition of focusing quadrupoles between the RFQ and inflector
The optimized spiral inflector, with $\mathrm{H}_{2}^{+}$ tagged through, is illustrated in Fig.3.17.
Figure 3.17: Illustrations of the optimized spiral inflector. Note the v-shape to the electrodes as well as the non-parallel angle seen in the bottom view, and the angled wedge visible in the top view.
@@
We thus consider two similar cases. Both have identical optimized spiral inflectors. The difference is whether or not a grounded collimator is placed at the entrance to further narrow the entering beam beyond what the quadrupoles have provided. This collimator is a circular aperture which, for this study, was chosen to be as large as possible while guaranteeing that less than one percent of $\mathrm{H}_{2}^{+}$ impact the electrodes. The resulting efficiencies and $\mathrm{H}_{2}^{+}$-electrode collisions are given in Table 3.3.
Table 3.3: Comparison of efficiency and ion-electrode impact for the optimized vs traditional spiral inflector.
@@
Traditional Inflector (Coll.) 39% $<1$%
Optimized Inflector (Coll.) 93% $<1$%
We also need to consider the impact of the new inflector on the quality of the beam which is successfully transmitted. These changes are characterized in Table 3.4. Detailed histograms showing the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ at the entrance and exit of the inflector are shown in Fig. 3.18.
Table 3.4: Comparison of transmitted beam parameters at the inflector exit.
$\sigma_{x}$ [mm] $\sigma_{y}$ [mm] $\sigma_{z}$ [mm] $\sigma_{r}$ [mm] $\sigma_{KE}$ [keV]
@@
Figure 3.18: Illustration of the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ prior to the inflector entrance vs at the inflector exit.
@@ 3.4.5 Orbit Matching and Beam Centering
Figure 3.19: Beam tracking result from the AIMA central region study. Left beam centering and acceleration on top of a view of the electrodes. Colors correspond to prticle phase. Right: 1 cm turn separation in the fourth turn with a single collimator. From [aima:central].
@@
A preliminary design study of the IsoDAR central region was performed by the company AIMA in France [aima:central]. The AIMA central region model is shown in Fig. 3.19. This central region was designed with a slightly different magnetic pole shape and with its own matched spiral inflector, using the ouput beam from an early RFQ beam tracking and single electrostatic quadrupole transverse refocusing element. It demonstrated that the RFQ-generated beam could be injected and vertically focused. By using a single collimator (removing $\approx 42$% of the beam), an edge-to-edge turn separation of 1 cm was achieved in the fourth turn (approximately 1 MeV/amu).
@@ 3.4.6 Low Energy Collimation
Following injection, there is significant beam spread in the first few turns of the cyclotron that we need to manage. We do this by placing a series of collimators around the beam core in the early turns. We must do all of our collimation in the first 7 turns of the primary acceleration, corresponding approximately to a 2 MeV/amu limit, to avoid activating our collimators. The placement of the collimators is shown in Figure  3.21 The reduction in RMS spread is shown in Fig  3.41. In total, we have 15 collimators that remove 36.1% of the beam. The improvement in beam shape is shown in  3.20. With time, the halo slowly redevelops but to a lesser extent than it would without collimation.
Figure 3.21: Placement of 15 collimators in first 7 turns
It is necessary to make the beam as circular as possible following the low energy collimators so that the beam is well-matched to “vortex motion.” Vortex motion stabilizes the beam and significantly reduces beam growth over time. It arises from the interaction between the self-field of the beam and the external magnetic field of the cyclotron. Vortex motion is discussed in depth in Ref [baumgarten:vortex2].
3.4.7 Risks and Mitigation
@@ 3.5 Vacuum Chamber
The vacuum enclosure comprises five parts: the upper and lower internal disk and 3 vacuum chambers. Fig 3.22 shows the vacuum enclosure in detail and the dimensions are in Table 3.5. The vacuum sealing is done using 2 metallic O-rings per interface (e.g., from Omniseal [PrecisionPolymerMetal]) except for the radial interface where Polymer O-rings could be used. The interface for the accessories will be done using a sleeve design and radial O-ring, described in Fig. 3.22.
Figure 3.22: Cyclotron vacuum chamber details
@@
Upper and lower internal disk 36 Ø4620*463
Upper vacuum chamber 2.3 Ø4452*483
Lower vacuum chamber - median plane 2.3 Ø4452*483
Lower vacuum chamber - internal disk 1.8 Ø4386*380
Table 3.5: Vacuum part properties
@@ 3.6 Coil configuration
The CDR anticipated a slightly shorter coil than we now anticipate using. Specifically, the coil height has been increased from 200 to 240 mm. Its radial width remains at 250 mm. The total current used in the calculations is 180,850 amp-turns in each of the half coils. The apparent current density is thus 3.01 A/mm2, i.e. $5\%$ lower than the 3.17 A/mm2 foreseen in the CDR. For now, no extraction channels are anticipated in the return yoke. If they are added, an increase in main coil current will be required
Figure 3.23: Left: Schematic view of a cross-section of the main coil, with six pancakes per half coil. Right: Possible configuration of the winding, using a square conductor with a round cooling channel in its center.
In Fig. 3.23 a schematic view of the coil is given. Also, a possible configuration of winding and cooling is calculated. Although this configuration does not exactly sum to a cross-section of 240 by 250 mm2, it shows that a cooling configuration is feasible.
@@ 3.7 Cyclotron RF System
Table 3.6: Details of the RF design.
@@
We consider these in the Sections below. The numerical parameters of the RF system are listed in Table 3.6.
@@ 3.8 RF Cavities (Dees, Stems and Liners)
The RF cavities of the cyclotron have been designed using CST microwave studio [cst:microwave_studio] in order to define the shape of the Dee and the Dee stems. We optimized our designs to reach the right resonance frequency, satisfy the Dee voltage Law, and minimize power dissipation. Ultimately, this improves the quality factor (Q-factor).
Figure 3.24: Section view of the RF cavity.
Figure 3.25: Outer view of the RF cavity.
The Dees and cavity walls are all made of oxygen-free high-conductivity (OFHC) copper, which has good electrical conductivity and good solderability. The Dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm. The four round stems support the Dee and provide the right inductance in order to produce the required resonance frequency. Their size and location produce the desired Dee voltage vs. radius profile. Each Dee will be equipped with vacuum pumping panels to efficiently pump the median plane. One Dee is equipped with one electrostatic deflector for extraction. This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the Dee. Figures 3.24 and 3.25 present the main elements composing the RF cavities.
@@ 3.8.1 Integration of preliminary central region in RF Cavities
In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA [aima:central] into the RF cavity modelling, several technical choices were made to ensure the effective implementation of this innovative configuration:
  • The central region design was trimmed at 150 mm from centre.
  • Both the Dee and counter dee original central region was removed by trimming the cavities at 250 mm from centre.
@@
The figure 3.26 shows the final integration of the central region. The lofted transitions can be better seen from top at figure 3.27. They leads to an acceptable linear vertical gap opening from 26 mm at a radius of 150 mm, to 50 mm at a radius of 250 mm. The section view of the model, presented at figure 3.28, shows the radial evolution of the vertical gap. In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil. Symmetry with regard to the median plane symmetry was maintained. Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
Figure 3.26: dee at central region close view
Figure 3.27: dee lofted transition top view
Figure 3.28: vertical gap evolution close view
@@ 3.8.2 RFQ Integration in RF Cavities
Some clearance provision has been set for the integration of the RFQ, as seen in figure 3.29
@@
Figure 3.29: Clearance for RFQ
@@ 3.8.3 RFQ getter pumps in RF Cavities
Incorporating getter pumps into the Dees allows us to improve our vacuum quality. By integrating three housing for SAES UHV1400 getter pumps [ModulesUHV] per dee, we aim to ensure highly efficient and reliable vacuum performance within the Cyclotron system. The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration. The figure 3.30 shows acceptable locations for three getter pump housings. Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity. The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
Figure 3.30: Provisional getter housings
@@ 3.8.4 RF Cavities design optimization
3.8.4.1 Performance requirement target
@@
The initial target value for the outer radius voltage was initially set at 250 kV, but according to our expertise, such high voltage will be very difficult to stand in cyclotron accelerator gaps of this size. It has been decided to reduce the target to 230 kV which remains a challenging value.
3.8.4.2 Parametric optimization of the design
We have intensively used parameters while building the CST model, to allow fast design exploration. We have retained four parameters to play with, for optimizing the cavity behavior. The parameters, shown at figure 3.31, are defined as a deviation from the early cavity design values.
@@ 3.8.4.2 Parametric optimization of the design
Figure 3.31: design optimization parameters
@@
Table 3.7: RF cavities optimization goals definition and achieved values.
@@
Mode1 frequency [MHz] 32.8 32.805
We selected the trust region algorithm and have defined scaled goals that would ensure convergence toward the required frequency and voltages. The trust region algorithm converged in a few tens of iterations. The definition of the optimization goals, and their target and reached values are contained in table 3.7. The table 3.8 shows the optimal values reached by the optimizer.
Table 3.8: RF cavities optimized design parameters.
Design parameter Value [mm]
@@ 3.8.5 RF cavity computation results
  • Frequency: The RF cavity operates at a frequency of 32.805 MHz, ensuring compatibility with the designed acceleration, and the RFQ integration,
  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.283, indicating that computed voltages at center would be 65.2 kV for a chosen voltage of 230 kV at the outer radius (2m),
  • CST computed quality factor (Q): 10586,
  • Estimated quality factor (Q): 7281,
  • Stored Energy (for max RF of 230 kV at 2 m): 4.06 Joules
  • CST Computed Losses: The computed losses per cavity using CST simulations are 79 kW,
@@
  • Maximal Surface Current: The rms surface current reaches a maximum of 17052.6 A/m in some reduced locations. Which leads to a maximum surface power of 43.5 W/cm². This maximal surface power is localized on the dee edges. Careful attention will be required for thermal management in these regions. For the rest of the surfaces, the thermal dissipation is well below 10 W/cm². See figure 3.36 and 3.37 While matching the requirement, these computation results provide valuable information for next steps of the Cyclotron’s RF cavity design including consideration in managing thermal effects.
Some illustrations of the electrical and magnetic fields computed by CST are given for reference at figures 3.32, 3.33, 3.34 and 3.35.
Figure 3.32: E field Vertical cross section
Figure 3.33: E-field Horizontal cross section
Figure 3.34: H-field horizontal cross section
Figure 3.35: H-field vertical cross section
Figure 3.36: Surface currents
3.8.5.1 Detailed power analysis
We will now detail the post treatment of the CST results that lead to the results presented in Sec. 3.8.5. The CST eigenmode solver results are computed for a given reference energy in the system (about 1J). They must then be scaled accordingly, to reach the expected voltage at a reference location. In our case, the reference location, is the orbit at 2m, and the desired voltage is 230 kV.
Figure 3.37: Surface current generating more than 10 W/cm2
The ratio of the desired voltage to the reference voltage at 2m is 2.013. We can scale the currents and voltages using this factor. We can also scale the energies, power and losses using this factor squared (formulae 3.1, 3.2 and 3.3). We thus have:
$E_{stored}=E_{stored,ref}\frac{V_{target}^{2}}{V_{ref}^{2}}$ (3.1)
$P_{loss}=P_{loss,ref}{V_{target}^{2}\over V_{ref}^{2}}$ (3.2)
$I{{}_{s}}=I_{s,ref}{V_{target}\over V_{ref}}$ (3.3)
Table 3.9: Optimized cavities, functional characteristics.
@@ 3.8.5.1 Detailed power analysis
Stored energy [J] 1 4.06
Voltage at on the 2m orbit kV] 114.2 230
Voltage at center [kV] 32.4 65.2
Maximal voltage at outer edge of the dee [kV] 124.6 250.9
Computed losses [kW] 19.48 78.9
Maximal surface current [A/m] 8470 17052.6
[…]
$R_{s}={1\over\sigma.\delta}$ (3.4)
$\delta=\sqrt{2\over{\sigma.\mu.\omega}}$ (3.5)
$P_{s}=R_{s}I_{s}^{2}$ (3.6)
Table 3.10: Cavity properties for computing surface dissipation
@@
Skin Depth $\delta$ $11.5$ µm
3.8.5.2 Radial voltage dependency
The voltages at different radii have been calculated by integrating the electrical field along circular arcs crossing the acceleration gaps, and reference curves crossing central gap and the dee outer edge gap. The transit time factor has not been considered. The computed voltages have then been scaled to match a target voltage of 230 kV at 2 m. The resulting absolute and normalized voltage radial dependency are given in figure 3.38 and table 3.11
Figure 3.38: Dee voltage radial dependency
Table 3.11: RF voltage radial dependency.
Radius [mm] Voltage [kV]

### v4

@@ 5.4.4 Characterization of Spiral Inflector
The result of adding optimized focusing quadrupoles on the beam distribution entering the spiral inflector is illustrated in Fig. 34.
Figure 34: Illustration of beam tightness at the entrance of the spiral inflector before and after the addition of focusing quadrupoles between the RFQ and inflector.
The optimized spiral inflector, with $\mathrm{H}_{2}^{+}$ tagged through, is illustrated in Fig. 35.
Figure 35: CAD renderings of the optimized spiral inflector. Note the v-shape to the electrodes as well as the non-parallel angle seen in the bottom view with the apertures removed, and the angled wedge visible in the top view.
@@
We thus consider two similar cases. Both have identical optimized spiral inflectors. The difference is whether or not a grounded collimator is placed at the entrance to further narrow the entering beam beyond what the quadrupoles have provided. This collimator is a circular aperture which, for this study, was chosen to be as large as possible while guaranteeing that less than one percent of $\mathrm{H}_{2}^{+}$ impact the electrodes. The resulting efficiencies and $\mathrm{H}_{2}^{+}$-electrode collisions are given in Table 8.
Table 8: Comparison of efficiency and ion-electrode impact for the optimized versus traditional spiral inflector.
@@
Traditional Inflector (Coll.) 39% <1<1%
Optimized Inflector (Coll.) 93% <1<1%
We also need to consider the impact of the new inflector on the quality of the beam which is successfully transmitted. These changes are characterized in Table 9. Detailed histograms showing the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ at the entrance and exit of the inflector are shown in Fig. 36.
Table 9: Comparison of transmitted beam parameters at the inflector exit.
$\sigma_{x}$ ($m$) $\sigma_{y}$ ($m$) $\sigma_{z}$ ($m$) $\sigma_{r}$ ($m$) $\sigma_{KE}$ (keV)
@@
Figure 36: Kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ before the inflector entrance versus at the inflector exit.
@@ 5.4.5 Orbit Matching and Beam Centering
Figure 37: Beam tracking result from the AIMA central region study. Left: beam centering and acceleration on top of a view of the electrodes. Colors correspond to prticle phase. Right: 1 cm turn separation in the fourth turn with a single collimator. From [64].
@@
A preliminary design study of the IsoDAR central region was performed by the company AIMA in France [64]. The AIMA central region model is shown in Fig. 37. This central region was designed with a slightly different magnetic pole shape and with its own matched spiral inflector, using the ouput beam from an early RFQ beam tracking and single electrostatic quadrupole transverse refocusing element. It demonstrated that the RFQ-generated beam could be injected and vertically focused. By using a single collimator (removing $\approx 42$% of the beam), an edge-to-edge turn separation of 1 cm was achieved in the fourth turn (approximately 1 MeV/amu).
@@ 5.4.6 Low Energy Collimation
Figure 38: Placement of 16 collimators in the first seven turns. All collimators are $10\times 20$ $m$.
Following the injection, there is a significant beam spread in the first few turns of the cyclotron that we need to manage. We do this by placing a series of collimators around the beam core in the early turns. We must do all of our collimation in the first 7 turns of the primary acceleration, corresponding approximately to a 2 MeV/amu limit, to avoid activating our collimators. The placement of the collimators is shown in Fig.  38. The reduction in RMS is shown in Fig. 59. In total, we have 16 collimators that remove 36.4% of the beam. The improvement in beam shape is shown in Fig. 39. With time, the halo slowly redevelops but to a lesser extent than it would without collimation.
Figure 39: Snapshots of the bunch in radial-longitudinal coordinates (aka local frame of the bunch) demonstrating the reduction of beam halo by collimators.
It is necessary to make the beam as circular as possible following the low energy collimators so that the beam is well-matched to “vortex motion.” Vortex motion stabilizes the beam and significantly reduces beam growth over time. It arises from the interaction between the self-field of the beam and the external magnetic field of the cyclotron. Vortex motion is discussed in depth in Ref. [5].
5.5 Risks and Mitigation
@@ 5.6 Vacuum Chamber
The vacuum enclosure comprises five parts: the upper and lower internal disk and 3 vacuum chambers. Fig. 40 shows the vacuum enclosure in detail and the dimensions are in Table 10. The vacuum sealing is done using 2 metallic O-rings per interface (e.g., from Omniseal [65]) except for the radial interface where Polymer O-rings could be used. The interface for the accessories will be done using a sleeve design and radial O-ring, described in Fig. 40.
Figure 40: Cross-section view of a CAD rendering of the HCHC-60 cyclotron, calling out vacuum chamber details.
Table 10: Vacuum part properties
@@
Upper and lower internal disk 36 ⌀4620*463
Upper vacuum chamber 2.3 ⌀4452*483
Lower vacuum chamber - median plane 2.3 ⌀4452*483
Lower vacuum chamber - internal disk 1.8 ⌀4386*380
@@ 5.7 Coil configuration
The CDR anticipated a slightly shorter coil than we now anticipate using. Specifically, the coil height has been increased from 200 to 240 mm. Its radial width remains at 250 mm. The total current used in the calculations is 180,850 amp-turns in each of the 180 half coils. The apparent current density is thus 3.01 A/mm2, i.e. $5\%$ lower than the 3.17 A/mm2 foreseen in the CDR. For now, no extraction channels are anticipated in the return yoke. If they are added, an increase in main coil current will be required
Figure 41: Left: Schematic view of a cross-section of the main coil, with six pancakes per half coil. The parameters of these coils are given in Table 11.
In Fig. 41 a schematic view of the coil is given. We show a possible configuration of winding and cooling in Table 11. Although this configuration does not exactly sum to a cross-section of 240 by 250 mm2, it shows that a cooling configuration is feasible.
@@
Table 11: Proposed coil configuration for a single pancake. The HCHC-60 will require 12 pancakes.
Parameter Value
# of axial turns 2
# of radial turns 12
Wire size $20.5\times 20.5$ mm2
Cooling hole diameter 15 mm
Pancake width 251 mm
Copper weight 708 kg
Current $1256\times 24$ Ampere-turns
Power 45 kW
Water flow 21.5 L/min
Pressure drop 12.3 kg/cm2
@@ 5.8 Cyclotron RF System
Table 12: Important parameters of the RF cavities. These are per cavity. The HCHC-60 will require four RF cavities.
@@
We consider these in the Sections below. The numerical parameters of the RF system are listed in Table 12.
@@ 5.9 RF Cavities (Dees, Stems and Liners)
The RF cavities of the cyclotron have been designed using CST microwave studio [66] in order to define the shape of the dee and the dee stems. We optimized our designs to reach the right resonance frequency, satisfy the dee voltage Law, and minimize power dissipation. Ultimately, this improves the quality factor (Q-factor).
Figure 42: Section view of the HCHC-60 double-stem, double-gap RF cavity, calling out the most important internal features.
Figure 43: CAD rendering of the HCHC-60 double-stem, double-gap RF cavity, calling out the most important external features.
The dees and cavity walls are all made of OFHC copper, which has good electrical conductivity and good solderability. The dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm. The four round stems support the dee and provide the right inductance in order to produce the required resonance frequency. Their size and location produce the desired Dee voltage versus radius profile. Each dee will be equipped with vacuum pumping panels to efficiently pump the median plane. One dee is equipped with one electrostatic deflector for extraction. This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the dee. Figures 42 and 43 present the main elements composing the RF cavities.
@@ 5.9.1 Integration of preliminary central region in RF Cavities
In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA [64] into the RF cavity modeling, several technical choices were made to ensure the effective implementation of this innovative configuration:
  • The central region design was trimmed at 150 mm from center.
  • Both the dee and dummy-dee original central region was removed by trimming the cavities at 250 mm from center.
@@
Figure 44 shows the final integration of the central region. The lofted transitions can be better seen from the top view in Fig. 45. They leads to an acceptable linear vertical gap opening from 26 mm at a radius of 150 mm, to 50 mm at a radius of 250 mm. The section view of the model, presented in Fig. 46, shows the radial evolution of the vertical gap. In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil. Symmetry with regard to the median plane symmetry was maintained. Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
Figure 44: CAD rendering of the dee–close-up of the central region, showing the spiral inflector housing and field-shaping posts.
Figure 45: CAD rendering of the dee–top view, showing the spiral inflector housing and field-shaping posts.
Figure 46: CAD rendering of the dee–side view, showing the radial evolution of the gap.
@@ 5.9.2 RFQ Integration in RF Cavities
Some clearance provision has been set for the integration of the RFQ, as seen in Fig. 47:
@@
Figure 47: CAD rendering of the dee, showing the clearance for RFQ.
@@ 5.9.3 RFQ getter pumps in RF Cavities
Incorporating getter pumps into the dees allows us to improve our vacuum quality. By integrating three housing for SAES UHV1400 getter pumps [67] per dee, we aim to ensure highly efficient and reliable vacuum performance within the cyclotron system. The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration. Figure 48 shows acceptable locations for three getter pump housings. Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity. The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
Figure 48: CAD rendering of the deer, showing the provisional getter housings (rectangular recesses).
@@ 5.9.4 RF Cavities design optimization
Performance requirement target
@@
The initial target value for the outer radius voltage was initially set at 250 kV, but from extensive prior experience we expect such high voltage will be very difficult to maintain in cyclotron accelerator gaps of this size. It has been decided to reduce the target to 230 kV which remains a challenging, albeit manageable, value.
Parametric optimization of the design
We have intensively used parameters while building the CST model, to allow fast design exploration. We have retained four parameters to play with, for optimizing the cavity behavior. The parameters, shown in Fig. 49, are defined as a deviation from the early cavity design values.
@@
Figure 49: CAD rendering of the dee, overlayed with an indication of the design optimization parameters.
@@
Table 13: RF cavity optimization goals definition and achieved values.
@@
Mode 1 frequency 32.8 MHz 32.805 MHz
Table 14: RF cavities optimized design parameters.
Design parameter Value (mm)
@@
We selected the trust region algorithm and have defined scaled goals that would ensure convergence toward the required frequency and voltages. The trust region algorithm converged in a few tens of iterations. The definition of the optimization goals, and their target and reached values are contained in Table 13. Table 14 shows the optimal values reached by the optimizer.
@@ 5.9.5 RF cavity computation results
  • Frequency: The RF cavity operates at a frequency of 32.805 MHz, ensuring compatibility with the designed acceleration and the RFQ integration.
  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.283, indicating that computed voltages at center would be 65.2 kV for a chosen voltage of 230 kV at the outer radius (2 m).
  • CST computed quality factor (Q): 10586
  • Estimated quality factor (Q): 7281
  • Stored Energy (for max RF of 230 kV at 2 m): 4.06 J
  • CST Computed Losses: The computed losses per cavity using CST simulations are 79 kW.
@@
  • Maximal Surface Current: The RMS surface current reaches a maximum of 17052.6 A/m in some reduced locations. Which leads to a maximum surface power of 43.5 W/cm2. This maximal surface power is localized on the dee edges. Careful attention will be required for thermal management in these regions. For the rest of the surfaces, the thermal dissipation is well below 10 W/cm2. See Figs. 54 and 55. While matching the requirement, these computation results provide valuable information for next steps of the cyclotron’s RF cavity design including consideration in managing thermal effects.
Some illustrations of the electrical and magnetic fields computed by CST are given for reference in Figs. 50, 51, 52 and 53.
Figure 50: RF cavity: E-field vertical cross section, calculated with CST.
Figure 51: RF cavity: E-field horizontal cross section, calculated with CST.
Figure 52: RF cavity: H-field horizontal cross section, calculated with CST.
Figure 53: RF cavity: H-field vertical cross section, calculated with CST.
Figure 54: RF cavity: Surface currents, calculated with CST.
Detailed power analysis
We will now detail the post treatment of the CST results that lead to the results presented in Section 5.9.5. The CST eigenmode solver results are computed for a given reference energy in the system (about 1 J). They must then be scaled accordingly, to reach the expected voltage at a reference location. In our case, the reference location, is the orbit at 2 m, and the desired voltage is 230 kV.
Figure 55: RF cavity: Surface current generating more than 10 W/cm2
The ratio of the desired voltage to the reference voltage at 2 m is 2.013. We can scale the currents and voltages using this factor. We can also scale the energies, power and losses using this factor squared (formulae 1, 2 and 3). We thus have:
$E_{stored}=E_{stored,ref}\frac{V_{target}^{2}}{V_{ref}^{2}}$ (1)
$P_{loss}=P_{loss,ref}{V_{target}^{2}\over V_{ref}^{2}}$ (2)
$I{{}_{s}}=I_{s,ref}{V_{target}\over V_{ref}}$ (3)
Here, $E_{stored}$ and $E_{stored,ref}$ are the actual stored anergy and the reference stored energy from the 1 J CST solution. $P_{loss}$ and $P_{loss,ref}$ are the actual power loss and the reference power loss from the 1 J CST solution. $V_{target}$ is the desired cavity accelerating voltage and $V_{ref}$ is the voltage obtained from the 1 J CST solution.
Table 15: Optimized cavities, functional characteristics.
@@
Stored energy (J) 1 4.06
Voltage at on the 2m orbit (kV) 114.2 230
Voltage at center (kV) 32.4 65.2
Maximal voltage at outer edge of the dee (kV) 124.6 250.9
Computed losses (kW) 19.48 78.9
Maximal surface current (A/m) 8470 17052.6
CST provides the maximal surface current $I_{s}$ in the cavity which isn’t a directly useful value, but we can compute from it the maximal surface power dissipation $P_{s}$, which will be useful for designing the cooling system sizing and layout. The physical properties needed for this determination are included in Table 16. We first need to compute the surface resistance $R_{s}$ whose formula, 4, uses the conductivity $\sigma$ and the RF skin depth $\delta$ (see formulae 4, 5 and 6). we end up with a maximal surface dissipation Power $43.5$ W/cm2. This value, despite being high, only concerns the very small surface of the dee edges, and should not be considered as an issue for the design of the cooling system.
$R_{s}={1\over\sigma\cdot\delta}$ (4)
$\delta=\sqrt{2\over{\sigma\cdot\mu\cdot\omega}}$ (5)
$P_{s}=R_{s}I_{s}^{2}$ (6)
Here, $\omega=2\pi f$ is the angular frequency and $\mu=\mu_{r}\cdot\mu_{0}$ the product of relative permeability and vacuum permeability.
Table 16: Cavity properties for computing surface dissipation
@@
[…]
Radial voltage dependency
The voltages at different radii have been calculated by integrating the electrical field along circular arcs crossing the acceleration gaps, and reference curves crossing central gap and the dee outer edge gap. The transit time factor has not been considered. The computed voltages have then been scaled to match a target voltage of 230 kV at 2 m. The resulting absolute and normalized voltage radial dependency are given in Fig. 56 and Table 17.
Figure 56: Radial dependency of the potential difference (“accelerating voltage”) across the accelerating gaps.
Table 17: RF voltage radial dependency.
Radius (mm) Voltage (kV)